OpenStudy (anonymous):

on each flight an airplane is filled with 10 quarts of oil. the airplane adds an additional 1/2qt of oil for each planned hour of flying. write a equation that can be used to determine the planned flying time (h) in hours for a airplane filled with 13qts of oil equation: that airplane flies 3300 miles. what is the average speed in miles per hour of the airplane?

OpenStudy (anonymous):

please help!!

OpenStudy (gw2011):

Let h=planned flying time in hours Let t=total quarts of oil filled in the plane Therefore, the equation can be set up as follows: (1/2)(h)=t-10 h=2(t-10) This is the equation Now substituting into the equation the known information you get: h=2(13-10) h=(2)(3) h=6 hours The average speed is calculated as follows: 3300 miles/6 hours=550 miles per hour